OpenStudy (anonymous):

how long will it take $10,000 invested at 5% compounded continuously to grow to $25,000?

OpenStudy (anonymous):

(10000) * (1.05^x) = 25000. Solve for x.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

from where you got 1.05?? the percentage is 0.05^x

hero (hero):

Continuous Compound Formula: \[A = Per^t\] A=Account Balance = 25,000 P=Principal = 10,000 r=rate (decimal form) = 0.05 t=time (in years) = what you need to find \[25,000 = 10,000e^{(0.05)t}\]

hero (hero):

In order to solve this, you will eventually, have to take logs. For now, start by dividing both sides by 10,000

hero (hero):

\[\frac{25000}{10000} = e^{(0.05)t}\]\[2.5 = e^{(0.05)t}\]

hero (hero):

Next, take logs of both sides: \[\log(2.5) = \log(e^{(0.05t)})\] \[log(2.5) = t\log(e^{0.05})\]

OpenStudy (anonymous):

i divided 25,000/10,000 = 10,000e(0.05)t/10,000 = 2.5=1.0512t both side divided by 1.0512 t=2.4

hero (hero):

\[\frac{\log(2.5)}{\log(1.051271)} = t\]

OpenStudy (anonymous):

ok

hero (hero):

We're using logs..so please remember that

OpenStudy (anonymous):

i punch it in the calculator and i got 18.33 for t

hero (hero):

Good job

OpenStudy (anonymous):

thanks to you

hero (hero):

If you use my formula, you get the most precise result.

hero (hero):

If you use virgil's formula, you still get 18, but you get the wrong approximation of months.

OpenStudy (anonymous):

yeah but i am really dumb when it come to logs and logarithm

hero (hero):

Well, there are some things you can do to make your life easier. My rule is to "only take logs when you need to".

OpenStudy (anonymous):

i will follow that too thank you